OpenStudy (keigh2015):

Choose the correct simplification of the expression (7x − 3)(4x2 − 3x − 6). 28x3 − 33x2 − 33x − 18 28x3 + 33x2 − 33x + 18 28x3 − 51x2 − 33x + 18 28x3 − 33x2 − 33x + 18

OpenStudy (keigh2015):

@iGreen @brina15

OpenStudy (anonymous):

I'm helping you now give me two seconds.

OpenStudy (igreen):

Just multiply the two expressions together.

OpenStudy (igreen):

\(\sf (7x - 3) \times (4x^2 − 3x − 6)\)

OpenStudy (keigh2015):

28?

OpenStudy (anonymous):

First work with the first term, 7x * 4x^2 and you get 28x^3. Then multiply the first term of the binomial and second term of the trinomial. so 7x * -3x what do you get? @Keigh2015

OpenStudy (igreen):

No, it's kind of like the last question.

OpenStudy (igreen):

Multiply all the terms in the first parenthesis to all the terms in the second parenthesis.

OpenStudy (anonymous):

@iGreen let me help her I'm showing her an easy way.

OpenStudy (igreen):

\(\sf 7x \times 4x^2\) \(\sf 7x \times -3x\) \(\sf 7x \times -6\) \(\sf -3 \times 4x^2\) \(\sf -3 \times -3x\) \(\sf -3 \times -6\)

OpenStudy (igreen):

Alright @alexistheking777

OpenStudy (anonymous):

@Keigh2015 You see what iGreen posted? Give me the answers to the problems he posted. It is part of the equation and will help you.

OpenStudy (keigh2015):

ok hold on

OpenStudy (igreen):

Yeah, after you find those just combine any like terms and you're done..

OpenStudy (anonymous):

Have you got your answers?

OpenStudy (keigh2015):

almost

OpenStudy (keigh2015):

In this order 28x^2 -21x -42x -12x^2 9x 18

OpenStudy (anonymous):

Your right for all the ones except the first two. The numbers are right but you forgot to combine the exponents so think of x^2 * x as x^2+1 since x has the exponential value of 1. So you would get an exponential value of 3. So it would be 28x^3. And same goes for x * x so it would be x^1+1 which is 2. So it should be -21x^2. 28x^3 -21x^2 -42x -12x^2 9x 18 Now combine like terms (that is terms with same exponent. So what is -21x^2 + (-12x^2)

OpenStudy (anonymous):

@Keigh2015

OpenStudy (keigh2015):

28x^3-33x^2

OpenStudy (anonymous):

Yes because 28x^3 is the only term with an exponential value of 3 and -21 + (-12) is like -21 - 12 which is 33. Then you replace the x and exponential value so its -33x^2. Good job, now we can eliminate answe C since the second term is 51x^2 which is wrong. So now what is -42x + 9x?

OpenStudy (keigh2015):

-33x

OpenStudy (anonymous):

By the way we can also eliminate B as well since the second term is positive and ours is negative. And yes it is -33x. Last step is to add the 18 to your equation so combine all those answers you just gave me.

OpenStudy (keigh2015):

28x^3-33x^2-33x+18

OpenStudy (anonymous):

Correct so when you look at the answers the only one that matches it is D. I hope this helped you don't forget to close this question and choose the person that was the greatest help to you by hitting best response and have a wonderful day. @Keigh2015

OpenStudy (keigh2015):

thank you for the help and you too @alexistheking777

OpenStudy (anonymous):

No problem.